OpenStudy (anonymous):

HEEEEEEEEEEEELPP!!!!!!!!!! Solve: 1 1/2n+3 1/2=14 not answer just way to work it out

OpenStudy (anonymous):

I said I can't understand what it says. =/ Use parenthesis please

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

1 and 1/2 n +3 and 1/2=14

OpenStudy (anonymous):

\[1+\frac{1}{2n+3}+\frac{1}{2}=14\] Is that it?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[1+\frac{1}{2}(n+3)+\frac{1}{2}=14\] This?

OpenStudy (anonymous):

im confused